# Structure-Informed Estimation for Pilot-Limited MIMO Channels via Tensor Decomposition

### v1

@@ 5.7 Discussion
The experimental results yield three main findings:
@@
    Tensor priors enable pilot-limited estimation. Low-rank Tucker completion improves NMSE by $\sim$2–12 dB over LS across the sweep, exceeding 10 dB for $\rho\geq 0.10$, and outperforms OMP at all pilot ratios. The empirical thresholds show that, at moderate-to-high SNR, pilot density is the dominant factor for achieving $\mathrm{NMSE}\leq 10^{-2}$.
@@
    Decomposition choice depends on channel structure. In the synthetic specular model, the CP baseline outperforms Tucker at moderate-to-high SNR (about 3–4 dB at 15–20 dB), while differences are small at low SNR and CP can be slightly worse. In DeepMIMO, Tucker is consistently slightly better than CP, indicating improved robustness under model mismatch. These results suggest CP for channels well approximated by sparse multipath, and Tucker for richer scattering conditions.
@@
    Hybrid modeling bridges the model–data gap. Tensor-NN yields 24–44% additional NMSE reduction on DeepMIMO relative to pure tensor methods and surpasses the ResNet baseline for $\rho\geq 0.04$, indicating that residual learning captures deviations beyond dominant low-rank structure.
@@ 5.8 Limitations
  • Rank selection and sensitivity. The tensor methods require multilinear ranks (or CP rank) to be specified a priori. In the current implementation, ranks are fixed per experiment and, in the sample-complexity study, are tied to the number of paths via $(R_{r},R_{t},R_{f})=(L,L,L{+}1)$. Performance can degrade under rank mismatch, and a principled, data-driven rank selection strategy remains necessary for robust deployment.
  • High-mobility and time-varying channels. The experiments assume quasi-static channels and estimators operate on single snapshots. Rapid mobility, Doppler spread, and temporal correlation are not modeled; extending the framework to spatio-temporal estimation or channel tracking is left for future work.
  • Neural training requirements. The neural components (Tensor-NN residual and ResNet baselines) are trained offline on paired $(\mathcal{H},\mathcal{Y}_{\Omega})$ data using multi-SNR and multi-pilot regimes. In deployment, this may require representative training data (or synthetic surrogates) and potential fine-tuning to mitigate distribution shift across scenarios, environments, or array configurations.
  • Synchronization and hardware impairments. The observation model includes only entry-wise sampling with additive Gaussian noise. Effects such as carrier-frequency offset, phase noise, IQ imbalance, nonlinear power amplifiers, low-resolution quantization, and other hardware imperfections are not modeled. Their impact on tensor completion and neural residual learning remains an open question.
@@ 5.9 Reproducibility
All experiments were conducted on an Apple M3 Pro laptop (11-core CPU, 14-core GPU, 18 GB unified memory) using Python 3.10 with PyTorch 2.1 and Metal Performance Shaders (MPS) acceleration. A fixed random seed (base_seed=42) ensures deterministic channel generation, pilot sampling, and network initialization across all experiments.
@@
DeepMIMO channels are generated using the official dataset API [34] with configuration files provided in the repository. Researchers can reproduce all figures and tables by executing the provided shell scripts with default parameters.
@@ 6 Conclusion
This paper presented a structure-informed framework for pilot-limited MIMO channel estimation that combines low-rank tensor completion with neural residual learning. Unlike prior tensor-based methods that assume fully observed received signal tensors, the proposed approach addresses the fundamentally underdetermined inverse problem arising from sparse pilot observations in fully digital MIMO-OFDM architectures.
The main contributions and findings are summarized as follows:
  1. 1.
    Tensor completion formulation: Modeling pilot-limited estimation as low-rank tensor completion yields $\sim$2–12 dB NMSE improvement over LS across the sweep (exceeding 10 dB for $\rho\geq 0.10$) and outperforms OMP at all pilot ratios, demonstrating that algebraic structure induced by sparse multipath propagation provides effective regularization under severe undersampling.
  2. 2.
    Decomposition selection: In the synthetic specular model, the CP baseline outperforms Tucker by about 3–4 dB at moderate-to-high SNR (15–20 dB), while differences are small at low SNR and CP can be slightly worse. In DeepMIMO, Tucker is consistently slightly better than CP, indicating greater robustness under model mismatch and richer scattering.
  3. 3.
    Hybrid tensor–neural architecture: Tensor-NN achieves 24–44% additional NMSE reduction on DeepMIMO ray-tracing channels relative to pure tensor methods and surpasses the ResNet baseline for $\rho\geq 0.04$, indicating that residual learning captures deviations beyond dominant low-rank structure.
  4. 4.
    Sample complexity: Recovery threshold analysis shows that the minimum number of pilot observations scales approximately as $|\Omega|_{\min}\approx C\cdot L(N_{r}+N_{t}+N_{f})$ rather than the ambient tensor size $N_{r}N_{t}N_{f}$, with empirical oversampling factors $C\approx 11$4141 exceeding CP-based theoretical minima, in part due to Tucker core parameters and algorithmic suboptimality.
The framework has limitations that warrant future investigation: (i) rank selection remains fixed per experiment rather than data-driven; (ii) high-mobility and time-varying channels are not modeled; (iii) the neural component requires offline training on representative data; and (iv) hardware impairments beyond additive noise are not considered.
Future directions include extension to time-varying channels via higher-order tensor decompositions incorporating temporal modes, reconfigurable intelligent surface (RIS)-aided systems with coupled tensor factors, adaptive rank selection strategies, and integrated sensing and communications (ISAC) scenarios requiring joint channel and target estimation. The proposed structure-informed approach naturally extends to emerging extremely large-scale MIMO (XL-MIMO) regimes, where complexity scaling with effective multilinear rank rather than ambient dimension becomes essential for computational tractability.

### v2

@@ 5.6 Discussion
The experimental results yield five main findings:
@@
    Tensor completion mitigates pilot scarcity. Fig. 4 shows that Tucker completion yields substantial NMSE gains over LS, LMMSE, and OMP across all pilot ratios. At $\rho=10\%$, Tucker achieves NMSE $\approx-11.25$ dB, corresponding to gains of $10.88$ dB over LS ($\approx-0.37$ dB) and $7.83$ dB over OMP ($\approx-3.42$ dB). LS remains below 1-1 dB across the full sweep, confirming that naive zero-filling cannot recover channel structure under sparse sampling. OMP improves over LS and LMMSE but still exhibits a performance floor around 3.5-3.5 dB due to dictionary mismatch and pilot scarcity. Notably, at the most extreme sparsity point ($\rho=2\%$), OMP is slightly better than Tucker by about $0.67$ dB, but Tucker becomes clearly dominant from $\rho\geq 4\%$ onward.
@@
    Decomposition performance depends on channel structure. Fig. 5 indicates that CP and Tucker perform similarly at low SNR (5-5 dB: CP $\approx-1.82$ dB versus Tucker $\approx-1.37$ dB, about $0.45$ dB difference), but CP is consistently better across the full SNR range under the synthetic specular model. The CP advantage increases with SNR and reaches $13.11$ dB at SNR = 20 dB (CP $\approx-25.00$ dB versus Tucker $\approx-11.89$ dB), consistent with the specular multipath structure of (7), where each propagation path contributes a rank-one component that matches the CP parameterization exactly.
    Importantly, CP maintains its advantage over Tucker even under the realistic DeepMIMO conditions of Experiment 3: for $\rho\geq 6\%$, CP outperforms Tucker by approximately 1122 dB across the pilot range. The exception occurs at extreme pilot scarcity ($\rho=2\%$), where CP exhibits heavy-tail divergence (mean NMSE $\approx+13.96$ dB) due to numerical instability of the WALS algorithm under severe undersampling, while Tucker’s alternating-projection scheme remains stable. This contrast highlights a key practical trade-off: CP offers superior reconstruction accuracy whenever sufficient pilots are available, but Tucker provides a safer fallback at the extreme undersampling boundary where CP initialization is unreliable. The hybrid Tensor–NN framework, which builds on Tucker completion, inherits this numerical stability while recovering performance through neural residual correction.
@@
    Hybrid Tensor–NN improves robustness on realistic channels. Fig. 6 (DeepMIMO, SNR = 10 dB) shows Tensor-NN as the best overall method across all pilot ratios. At $\rho=8\%$, Tensor-NN reaches 13.72-13.72 dB, improving by $2.26$ dB over CP ($\approx-11.46$ dB), $4.80$ dB over Tucker ($\approx-8.93$ dB), and $5.85$ dB over LMMSE ($\approx-7.87$ dB). At $\rho=20\%$, Tensor-NN remains best ($\approx-18.43$ dB), with a $1.00$ dB margin over CP ($\approx-17.43$ dB). CP exhibits severe heavy-tail instability at $\rho=2\%$ (mean NMSE $\approx+13.96$ dB), while Tensor-NN remains stable throughout.
  4. 4.
    Algebraic structure consistently outperforms unconstrained deep learning. The comparison with the ResNet baseline in Fig. 6 shows that Tensor-NN outperforms ResNet at every pilot density tested, with a monotonically growing margin: $1.53$ dB at $\rho=2\%$, $3.64$ dB at $\rho=4\%$, $3.58$ dB at $\rho=6\%$, $3.67$ dB at $\rho=8\%$, $3.91$ dB at $\rho=10\%$, $4.91$ dB at $\rho=15\%$, and $5.67$ dB at $\rho=20\%$. The global SNR reference of (16) ensures that the advantage of algebraic structure is not an artefact of the measurement setting. The monotonically growing gap indicates that structure-informed estimation scales more efficiently with pilot density than unconstrained learning, which is ultimately bounded by training data diversity rather than intrinsic channel geometry. These results demonstrate that the algebraic prior is the dominant source of performance gain across the full pilot-density range, and that the neural residual stage provides complementary correction for model mismatch rather than compensating for insufficient structure.
  5. 5.
    Sample complexity is governed by intrinsic channel structure. Experiment 4 shows that the minimum pilot ratio $\rho_{\min}$ and observation count $|\Omega|_{\min}$ increase monotonically with the number of propagation paths $L$. At SNR = 10 dB, no configuration reaches $\mathrm{NMSE}\leq 10^{-2}$ within the tested range, indicating a noise-limited regime where increasing pilot density alone cannot compensate for measurement noise. For SNR $\in$$\{20,\,30\}$ dB, recovery thresholds lie in the 12–25% pilot range and are primarily governed by $L$ rather than by the ambient tensor dimension $N_{r}N_{t}N_{f}$, consistent with Table 3.
@@ 5.7 Limitations
  • Rank selection and sensitivity. The tensor methods require multilinear ranks (or CP rank) specified a priori. The current implementation fixes ranks per experiment and, in the sample-complexity study, ties them to the number of paths via $(R_{r},R_{t},R_{f})=(L,L,L{+}1)$. Performance degrades under rank mismatch, and a principled, data-driven rank selection strategy remains necessary for robust deployment.
  • High-mobility and time-varying channels. The experiments assume quasi-static channels and single-snapshot estimators. Rapid mobility, Doppler spread, and temporal correlation fall outside the model; extending the framework to spatio-temporal estimation or channel tracking remains future work.
  • Neural training requirements. The neural components (Tensor-NN residual and ResNet baselines) train offline on paired $(\mathcal{H},\mathcal{Y}_{\Omega})$ data under multi-SNR and multi-pilot regimes. Deployment may require representative training data (or synthetic surrogates) and fine-tuning to mitigate distribution shift across scenarios, environments, or array configurations.
  • Synchronization and hardware impairments. The observation model covers only entry-wise sampling with additive Gaussian noise. Effects such as carrier-frequency offset, phase noise, IQ imbalance, nonlinear power amplifiers, low-resolution quantization, and other hardware imperfections fall outside the model. Their impact on tensor completion and neural residual learning remains an open question.
@@ 5.8 Reproducibility
All experiments ran on a Lenovo laptop (Intel Core i7, 16 GB RAM, NVIDIA GeForce RTX 2060 6 GB GPU) under Python 3.10 with PyTorch 2.1. A fixed random seed (base_seed=42) ensures deterministic channel generation, pilot sampling, and network initialization across all experiments.
@@
The experiments generate DeepMIMO channels using the official dataset API [34] with configuration files provided in the repository. Researchers can reproduce all figures and tables by executing the provided shell scripts with default parameters.
@@ 6 Conclusion
This paper proposed a structure-informed framework for pilot-limited wideband MIMO channel estimation combining low-rank tensor completion with neural residual learning. The formulation directly addresses the ill-posed inverse problem induced by sparse pilot observations by exploiting algebraic priors that arise universally from sparse multipath propagation and array geometry, without requiring site-specific environmental models. SNR is defined globally with respect to the average power per channel coefficient over the full tensor—independent of the pilot set $\Omega$ and pilot ratio $\rho$—ensuring strict comparability across all experiments and pilot densities.
Synthetic-channel experiments demonstrate that Tucker completion achieves $10.88$ dB NMSE improvement over LS and $7.83$ dB over OMP at $\rho=10\%$ pilot density. CP outperforms Tucker across the full SNR range under the specular multipath model, reaching a $13.11$ dB advantage at SNR = 20 dB, consistent with the rank-one multipath structure that matches the CP parameterization exactly. On DeepMIMO ray-tracing channels, CP maintains its advantage over Tucker for $\rho\geq 6\%$ but exhibits severe heavy-tail divergence at $\rho=2\%$ (mean NMSE $\approx+13.96$ dB), where Tucker’s alternating-projection scheme remains numerically stable. The proposed Tensor–NN estimator achieves $2.26$ dB gain over CP and $4.80$ dB over Tucker at $\rho=8\%$, and outperforms the purely data-driven ResNet baseline at every pilot density tested, with a monotonically growing margin from $1.53$ dB at $\rho=2\%$ to $5.67$ dB at $\rho=20\%$. This result confirms that algebraic structure is the dominant source of performance gain across the full pilot-density range, and that the neural residual stage provides complementary correction for model mismatch rather than substituting for structural priors.
[…]
These findings establish tensor priors as physically grounded, portable, and theoretically motivated constraints for pilot-limited estimation, while neural residual learning provides a principled mechanism to compensate model mismatch under realistic propagation conditions. Future work will address adaptive rank selection via model-order estimation, extensions to time-varying channels through joint space–frequency–time tensor modeling, integration with hybrid analog-digital architectures, and robustness analysis under hardware impairments including phase noise, nonlinearities, and low-resolution quantization.